OpenStudy (anonymous):

Three party-goers are in the corner of the ballroom having an intense argument. You walk over to settle the debate. They are discussing a function g(x). You take out your notepad and jot down their statements. Professor McCoy: She says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. Mr. Romano: He says that 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0. Correct the reasoning of any inaccurate reasoning by the party-goers in full and complete sentence.Make sure you reference any theorems that support your justifications.

terenzreignz (terenzreignz):

Interesting party, what with all these nerds >:) ...well, some nerds they are XD What are your thoughts so far? Any ideas?

OpenStudy (anonymous):

Haha well actually I don't get how to do this I'm so lost

terenzreignz (terenzreignz):

Okay, consider the Professor's theory first. If (x+2) is a factor of a function, does it automatically mean that 2 is a zero of the function?

OpenStudy (anonymous):

Yes?

terenzreignz (terenzreignz):

It does, does it? Consider h(x) = (x+2)(x+3) And let x = 2. what is h(x)?

OpenStudy (anonymous):

X+5x+6

terenzreignz (terenzreignz):

Yes, but x = 2, and what happens?

OpenStudy (anonymous):

You put 2 in every x so it would look like 2+10+6

terenzreignz (terenzreignz):

Is that zero? :P

terenzreignz (terenzreignz):

2 + 10 + 6 is that zero?

OpenStudy (anonymous):

No

terenzreignz (terenzreignz):

No. So 2 wasn't a zero of h(x)... was it? :P

OpenStudy (anonymous):

2+10+6

terenzreignz (terenzreignz):

18. Not zero. So 2 was not a zero of h(x), since h(2) is not zero. Understood?

OpenStudy (anonymous):

Yes !

terenzreignz (terenzreignz):

So, the professor is wrong...

OpenStudy (anonymous):

Ok so what about ms guerra

terenzreignz (terenzreignz):

Ms. Guerra gave the precise definition of a zero of a function. Can't argue with her :P

OpenStudy (anonymous):

What about mr Romano?

terenzreignz (terenzreignz):

Ever synthetic division?

OpenStudy (anonymous):

We didn't use synthetic division

terenzreignz (terenzreignz):

That's a problem. Synthetic division with r gives you the result when a polynomial is divided by (x-r)

OpenStudy (anonymous):

Ok well I know how to somewhat do that

OpenStudy (anonymous):

So do I divide (x+2) by something?

terenzreignz (terenzreignz):

No, all I mean is that synthetic division with 2 means it was divided by (x-2) and there was no remainder, right?

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

So, that means g(x) = (x-2)f(x) for some function f(x)

OpenStudy (anonymous):

Ok that it?

terenzreignz (terenzreignz):

Well, try having x=2 now.

OpenStudy (anonymous):

If I did this I got 1+3

terenzreignz (terenzreignz):

How?

OpenStudy (anonymous):

Ok we'll if I understood this right I plugged in my f(x) equation you showed me and I divided it by (x+2) did I do it wrong? Sorry I'm such a pain...

terenzreignz (terenzreignz):

You're not a pain, lol But in g(x) = (x-2)f(x) Just replace all the x's with 2. Show me what happens...step-by-step if you must ^_^

OpenStudy (anonymous):

Ok plug in 2 in all x's I got 0 f(2)

terenzreignz (terenzreignz):

0 times f(2) 0 times anything is 0 So you got g(2) equals zero after all.

OpenStudy (anonymous):

Ooo ok now I got what you where saying

terenzreignz (terenzreignz):

So, by the third nerd's reasoning, yes, it is a valid argument. Meaning it's only (ironically) the professor who's off. Got it? good :P

OpenStudy (anonymous):

Than you so much for helping me I have you a medal ;)

terenzreignz (terenzreignz):

No problem. Just polish it up and you'll be in shape ^_^

OpenStudy (anonymous):

gave* sorry and um thank you again.